OpenStudy (anonymous):

Write the equation in slope-intercept form. 2x + 3y = 18 A. y = 18 - 2x B. y = -2/3 x + 6 C. y = 3/2 x + 18 D. y = 18 - 3x

OpenStudy (mysesshou):

solve for y:

OpenStudy (mysesshou):

2x + 3y = 18 subtract 2x each side: 3y = 18 - 2x divide both sides by 3: y = 18/3 - 2/3x y = 6 - 2/3x rearrange: y = - 2/3x + 6